# Homogeneous Tokenizer Matters: Homogeneous Visual Tokenizer for Remote Sensing Image Understanding

### v1

@@ 4.2 Object Perception Module
HOOK treats semantically independent regions as the basic elements of vision. How do we find SIRs in an image? Gaining inspiration from a similar problem in NL tokenization—how to find semantically meaningful subwords—we examine an elegant solution provided by BPE[45]: first, we split the language into the smallest elements, such as splitting based on letters. Second, we calculate the most frequently occurring letter combinations in the corpus and construct new tokens based on those combinations, iterating this process. This approach corresponds to the ”Splitting and Merging” routing mentioned in Section 3.3. Our designed OPM consists of three main steps: 1. Splitting the image into several fine-grained seeds; 2. Merging the seeds into SIRs; 3. Stopping and merging.
@@
One of the most intuitive approaches inspired by BPE is to treat each pixel in the image as a seed and then to gradually merge them to form SIRs. However, the cost of pixels as initial seeds is too expensive, as a 224 $\times$ 224 image can have as many as 50,176 pixels. To address this issue, we introduce a hypothesis specifically for remote sensing images:
@@
In remote sensing images, there are no complete objects within a 4x4 pixel window.
Under Assumption 1, the size of any SIR is larger than 4×4, so dividing the image into several 4 $\times$ 4 pixel-sized seeds actually converts from ”Multiple Tokens Multiple Objects” to ”Same Object Multiple Tokens”, as shown in Figure 2 from square 4 to square 2.
In the implementation, we use convolutional modules to obtain the seeds, as described above. Specifically, we stack 2 layers of convolutional layers with a kernel size of 2 and a stride of 2 to extract a local feature vector from the 4 $\times$ 4 pixel window, which becomes the obtained seed. Additionally, to enhance the representation capability of the seeds and support the expansion of subsequent SIRs, we adopt the traditional CNN architecture and intersperse convolutional layers with a kernel size of 3, a stride of 1, a padding of 1, batch normalization layers, and rectified linear unit (ReLU) activation layers within the convolutional block. At the end of the convolutional block, we add a dimension projection layer with a kernel size of 1 and a stride of 1 to project the image features to a 512-dimensional space.
@@
How can seeds be expanded into SIRs? In natural language processing, the self-attention mechanism essentially relates and aggregates tokens, making tokens with semantic relationships more similar. This phenomenon has also been observed in various visual data[42, 47, 48, 49]. Inspired by this finding, we attempt to introduce a self-attention layer to make the seeds belonging to the same SIR more similar to achieve expansion from seeds to SIRs.
In the implementation, we adopted the SAM[50, 51] and did not use the standard self-attention module. Instead, we introduce local attention and global attention. First, local attention is used to establish relationships among local seeds, and then global attention is used to expand this relationship to a global scale. The ablation experiments in Section 5.5 demonstrate that local attention, global attention, and their stacking play important roles in the overall performance of HOOK.
@@
According to the analysis in Section 3.1, the ideal granularity of semantically independent regions depends on the specific image and task. Therefore, in the stop strategy for the continuous merging of seeds into SIRs, we did not introduce additional constraints or guidance. The expansion process of seeds naturally stops after passing through one layer of local attention and one layer of global attention, and it is jointly optimized with the entire model under the supervision of the specific task loss function.
Figure 4: HOOK is capable of adapting to tasks with different granularities. For sparse tasks represented by classification, the model can average the tokens output by the model and then pass them through a linear classification layer to output the classification results. For dense tasks represented by segmentation, the model can utilize the intermediate variable, the attention map from the OVM, to restore the number of tokens and then pass them to the segmentation head to output the segmentation results.
@@ 4.3 Object Vectorization Module
The OPM merges seeds into SIRs, where seeds belonging to the same SIR will be more similar. The OVM is aimed at the vectorization of SIRs into visual tokens, and the number of visual tokens can be arbitrarily adjusted to meet adaptability requirements.
A. Jaegle et al. proposed a perception model framework called the Perceiver[52], which consists of only cross-attention and self-attention layers. This model can perceive multiple modalities without changing its structure. The core idea is that the modality data serve as the key and value for cross-attention, while the model’s predefined latent variables serve as queries. The latent variables are repeatedly exposed to the original modality data through stacked cross-attention, continuously refining the model’s perceptual results. Moreover, because only the latent variables pass through self-attention, the model’s efficiency is greatly improved.
The success of the Perceiver has inspired us in two ways: (1) The predefined query in cross-attention has the ability to perceive semantic information in the key and value. (2) The number of tokens obtained through cross-attention is determined only by the query. Furthermore, from the calculation principle of cross-attention, as indicated by Equation 5, it can be observed that the operation of calculating similarity between the query and the key matches the way the OPM finds SIRs (i.e., making seeds belonging to the same SIR more similar).
$v^{\prime}=Softmax(\frac{QK^{T}}{\sqrt{d}})V$ (5)
Therefore, we implement the object vectorization module based on cross-attention. Specifically, we define $N$ learnable vectors $q$ as the query for cross-attention, while the seeds outputted by the object perception module serve as key and value. During the forward process, each vector in $q$ queries the seeds to retrieve the most similar seed set and aggregates them with weighting to form visual tokens.
In the specific implementation, we empirically found that when the dimension of cross-attention is defined as half of the object perception module dimension, i.e., 256, HOOK performs the best. Additionally, to align the dimensions of the visual tokens with the subsequent backbone network dimensions, we added an MLP after the cross-attention layer. The dimensions of the MLP and the learnable vector q are set to the dimensions of the backbone network. For example, if the backbone network is a standard ViT, the dimension is 768.
@@ 4.4 HOOK is image-agnostic and task-agnostic
HOOK is independent of specific images and tasks and is image-agnostic and task-agnostic.
First, HOOK can accept images of any size. HOOK uses a convolutional block to extract local features from the image (as detailed in Section 4.2). Since convolutional operations are not sensitive to image size, HOOK can handle images of any size. Considering that a large-scene image may lead to too many seeds in a 4 $\times$ 4 pixel window, which would increase the computational cost of the two self-attention layers in the OPM, we implemented adjustable seed sizes. In the experiments in Section 5.2, we demonstrated the changes in efficiency and accuracy due to variations in the convolution window size. We showed that even for large-scene remote sensing images, HOOK maintains high efficiency and accuracy.
Furthermore, HOOK demonstrates excellent performance on both dense tasks and sparse tasks. In sparse prediction tasks such as classification, the model only needs to output sparse discriminative information for the original image. As shown in Figure 4, after the tokens outputted by HOOK go through the backbone network, they can directly pass through a linear classification layer to produce classification results. In dense prediction tasks such as semantic segmentation, the model needs to output not only discriminative information but also pixel-level spatial information. We observed that the cross-attention in the OVM can output not only visual tokens but also intermediate variables: attention maps, which indicate the region represented by each visual token (i.e., region of the token mentioned in Section 3.2). Therefore, we use the attention map to reconstruct the original number of visual tokens, thus restoring spatial information. The experiments in Section 5.2 demonstrate that HOOK requires only 8 tokens to perform the semantic segmentation task.
5 Experiment
@@
A complete transformer-based visual model should consist of three main components: a visual tokenizer, a transformer backbone, and a task head. The visual tokenizer is responsible for converting raw images into token sequences, with the most popular visual tokenizer currently being Patch Embed. The transformer backbone serves as the core of the model, extracting and understanding the semantic information within the token sequences. The task head aligns the output from the backbone with specific downstream tasks. In this paper, we propose a new visual tokenizer called HOOK, which is aimed at replacing Patch Embed to enhance the model’s ability to understand visual data.
@@
To fairly and effectively test the performance of HOOK, unless specified otherwise, we used the same transformer backbone (12-layer transformer encoder) and task head (linear classifier for classification tasks & SegFormer[53] head for segmentation tasks) in all experiments. Notably, many improvements to visual tokenizers not only change the structure of the visual tokenizer but also introduce new modules or mechanisms into the backbone or task head[26, 28, 36, 40, 54]. To accurately assess the impact of the visual tokenizer on the model’s performance, our baseline for comparison excludes these types of methods and only includes work that focuses solely on improving the visual tokenizer.
@@
For the credibility of our experimental results, our first principle in selecting datasets is that they are publicly available and widely researched. Additionally, we aimed to validate that HOOK is image-agnostic and task-agnostic while testing its performance. Therefore, we tested the performance of HOOK on both sparse tasks, represented by classification, and dense tasks, represented by segmentation. Specifically, for the classification tasks, we chose the NWPU-RESISC45[18] and WHU-RS19 datasets[19, 20], with both datasets having fixed image sizes of 224 $\times$ 224. For the segmentation task, we selected the GID5 dataset[21], which has a fixed image size of 512 $\times$ 512.
The experiments demonstrated that HOOK possesses two fundamental properties—homogeneity and adaptability—and exhibits excellent performance and efficiency. Section 5.2 presents the test results of HOOK on the aforementioned two types of tasks and quantitatively compares the efficiency of HOOK with Patch Embed. The visualization results of the homogeneity of visual tokens are shown in Section 5.3. Additionally, we conducted more in-depth analysis and validation of the phenomena observed in the experiments. Section 5.4 experimentally verifies the simple intuition that under a stronger visual tokenizer, deeper backbone networks are redundant. The ablation results and analysis of the two types of attention mechanisms within the OVM are presented in Section 5.5, which shows that both attention mechanisms play important roles in improving the homogeneity of visual tokens, thereby affecting the final accuracy performance of the model. In Section 5.6, we discuss some limitations of HOOK to provide a comprehensive view of the HOOK method.
@@ 5.2 Main Results
We tested the performance of the model with HOOK as the visual tokenizer on two categories of tasks: sparse tasks and dense tasks. The essential difference between the two types of tasks is that sparse tasks only require sparse discriminative information, such as categories, while dense tasks require not only discriminative information but also dense spatial information. Representative sparse tasks include classification, re-identification, sentiment analysis, etc., while representative dense tasks include segmentation, detection, depth estimation, etc.
Table 1: Sparse and Dense Task with HOOK
Tokenizer Sparse Task (Classification) Dense Task (Segmentation)
Num of Tokens NWPU2:8(Acc1) WHU-RS19(Acc1) Num of Tokens GID5(mIoU)
Patch Embed 196 70.30 78.10 1024 67.79
Quadtree 100 70.99 81.05 400 49.18
VT 8 71.95 80.07 8 58.12
PnP-DETR 54 73.53 81.37 158 62.12
Conv-VGG19 49 74.46 83.66 256 53.73
Conv-ResNet50 49 75.17 81.37 256 68.81
HOOK(ours) 6 77.38 87.58 8 78.81
@@
Table 1 displays the experimental results. To accurately evaluate the impact of different visual tokenizers on the overall performance of the model, the baseline for comparison includes only works that optimize the visual tokenizer while keeping the transformer backbone and task head consistent. In our selected baseline, Qurdtree[35] and PnP-DETR[33] are patch-based methods. The former assumes that all patches in the image are unreasonable and uses GradCAM to identify significant regions, implementing the generation of patches with different granularities using quadtrees. The latter reveals that the number of patches is redundant, using a scoring mechanism to identify background and foreground regions and merging features of background areas. VT[42] is an object-oriented method that uses a simple convolution operation to assign each pixel in the image to one of several semantic groups, which are eventually mapped to a visual token. Additionally, some studies[25, 33] have suggested that the combination of convolution and a transformer leads to better performance for transformer-based models, so we also tested the performance of two classic convolutional networks (VGG[55] and ResNet[56]) as visual tokenizers.
A. Sparse Task

### v2

@@ 4.2 Object Perception Module
HOOK treats semantically independent regions as the basic elements of vision. How do we find SIRs in an image? To gain inspiration from a similar problem in NL tokenization, namely, finding semantically meaningful subwords, we examine an elegant solution provided by Byte-Pair Encoding (BPE)[47]: First, we split the language into the smallest elements, such as by splitting on the basis of letters. Second, we calculate the most frequently occurring letter combinations in the corpus and construct new tokens on the basis of those combinations, iterating this process. This approach corresponds to the ”splitting and merging” route mentioned in Section 3.3. The OPM consists of three main steps: 1. splitting the image into several fine-grained seeds; 2. merging the seeds into SIRs; and 3. stopping and merging.
@@
One of the most intuitive approaches inspired by Byte-Pair Encoding (BPE) is to treat each pixel in the image as a seed and then to gradually merge them to form SIRs. However, the cost of pixels as initial seeds is too high, as a 224 $\times$ 224 image can have as many as 50,176 pixels. To address this issue, we introduce a hypothesis specifically for remote sensing images:
@@
In remote sensing images, there are no complete objects within a 4$\times$4 pixel window.
Under Assumption 1, the size of any SIR is larger than 4$\times$4, so dividing the image into several 4$\times$4 pixel-sized seeds actually converts the scenario from ”multiple tokens multiple objects” to ”same object multiple tokens”, as shown in Figure 2 from square 4 to square 2.
In the implementation, we use convolutional modules to obtain the seeds, as described above. Specifically, we stack 2 convolutional layers with a kernel size of 2 and a stride of 2 to extract a local feature vector from the 4 $\times$ 4-pixel window, which becomes the obtained seed. Additionally, to enhance the representation capability of the seeds and support the expansion of subsequent SIRs, we adopt the traditional CNN architecture and intersperse convolutional layers with a kernel size of 3, a stride of 1, and a padding of 1 with batch normalization layers and rectified linear unit (ReLU) activation layers within the convolutional block. At the end of the convolutional block, we add a dimension projection layer with a kernel size of 1 and a stride of 1 to project the image features to a 512-dimensional space.
@@
How can seeds be expanded into SIRs? In natural language processing, a self-attention mechanism essentially relates and aggregates tokens, making tokens with semantic relationships more similar. This phenomenon has also been observed in various visual data[44, 49, 50, 51]. Inspired by this finding, we attempt to introduce a self-attention layer to make the seeds belonging to the same SIR more similar to achieve expansion from seeds to SIRs.
Considering that the token sequence from step 1 is much longer than that of Patch Embed and that multiple works[52, 53, 54, 55] have reported that self-attention suffers from decreased efficiency and attention degradation when processing long sequences, we adopt the SAM[55, 56] instead of the standard self-attention module.
Specifically, we introduce local attention and global attention into the SAM. First, local attention is utilized to establish relationships among seeds within a local window, which is typically constrained to a quarter of the image. Second, global attention expands these relationships to a global scale, functioning similarly to a standard self-attention layer. The ablation experiments in Section 5.7 demonstrate that local attention, global attention, and their stacking all play important roles in the overall performance of HOOK.
@@
According to the analysis in Section 3.1, the ideal granularity of semantically independent regions depends on the specific image and task. Therefore, in the stopping strategy for the continuous merging of seeds into SIRs, we did not introduce additional constraints or guidance. The expansion process of seeds naturally stops after passing through one layer of local attention and one layer of global attention, and it is jointly optimized with the entire model under the supervision of a specific task loss function.
Figure 4: HOOK is capable of adapting to tasks with different granularities. For sparse tasks represented by classification, the model can average the tokens and then pass them through a linear classification layer to output the classification results. For dense tasks represented by segmentation, the model can utilize the intermediate variable, namely, the attention map from the OVM, to restore the number of tokens and then pass them to the segmentation head to output the segmentation results.
@@ 4.3 Object Vectorization Module
The OPM merges seeds into SIRs, where seeds belonging to the same SIR will be more similar. The OVM is aimed at vectorizing SIRs into visual tokens, and the number of visual tokens can be arbitrarily adjusted to meet adaptability requirements.
A. Jaegle et al. proposed a perception model framework called the Perceiver[57], which consists of only cross-attention and self-attention layers. This model can perceive multiple modalities without changing its structure. The core idea is that the modality data serve as the keys and values for cross-attention, whereas the model’s predefined latent variables serve as queries. The latent variables are repeatedly exposed to the original modality data through stacked cross-attention, continuously refining the model’s perceptual results. Moreover, because only the latent variables pass through the self-attention layer, the model’s efficiency is greatly improved.
The success of the Perceiver inspires us in two ways: (1) The predefined query in cross-attention has the ability to perceive semantic information in the key and value. (2) The number of tokens obtained through cross-attention is determined only by the query.
Furthermore, from the calculation principle of cross-attention, as indicated by Equation 5, where $Q$ denotes predefined queries, $K$ and $V$ denote outputs of the OPM, $\sigma$ denotes the Softmax operation, and $d$ denotes the dimension of $Q$, the operation of calculating the similarity between the query and the key matches the way the OPM finds SIRs (i.e., making seeds belonging to the same SIR more similar).
$v^{\prime}=\sigma(\frac{QK^{T}}{\sqrt{d}})V$ (5)
Therefore, we implement the object vectorization module on the basis of cross-attention. Specifically, we define $N$ learnable vectors $q$ as the queries for cross-attention, while the seeds output by the object perception module serve as the keys and values. During the forward process, each vector in $q$ queries the seeds to retrieve the most similar seed set and aggregates them with weights to form visual tokens.
In the specific implementation, we empirically find that when the dimension of cross-attention is defined as half of the object perception module’s dimension, i.e., 256, HOOK performs the best. Additionally, to align the dimensions of the visual tokens with the subsequent backbone network dimensions, we add an MLP after the cross-attention layer. The dimensions of the MLP and the learnable vectors q are set to the dimensions of the backbone network. For example, if the backbone network is a standard ViT, the dimension is 768.
@@ 4.4 HOOK is image-agnostic and task-agnostic
HOOK is independent of specific images and tasks and is both image-agnostic and task-agnostic.
First, HOOK can accept images of any size. HOOK uses a convolutional block to extract local features from an image (as detailed in Section 4.2). Since convolutional operations are not sensitive to image size, HOOK can handle images of any size. Considering that a large-scene image may lead to too many seeds in a 4$\times$4 pixel window, which would increase the computational cost of the two self-attention layers in the OPM, we implement adjustable seed sizes. In the experiments in Section 5.2, we demonstrated the changes in efficiency and accuracy due to variations in the convolution window size. We showed that even for large-scale remote sensing images, HOOK maintains high efficiency and accuracy.
Furthermore, HOOK demonstrates excellent performance on both dense tasks and sparse tasks. In sparse prediction tasks such as classification, the model needs to output only sparse discriminative information for the original image. As shown in Figure 4, after the tokens output by HOOK pass through the backbone network, they can directly pass through a linear classification layer to produce classification results. In dense prediction tasks such as semantic segmentation, the model needs to output not only discriminative information but also pixel-level spatial information. We observe that the cross-attention in the OVM can output not only visual tokens but also intermediate variables, namely, attention maps that indicate the region represented by each visual token (i.e., the region of the token mentioned in Section 3.2). Therefore, we use the attention map to reconstruct the original number of visual tokens, thus restoring spatial information. The experiments in Section 5.2 demonstrate that HOOK requires only 8 tokens to perform the semantic segmentation task.
5 Experiments
The experiments demonstrate that HOOK possesses two fundamental properties, namely, homogeneity and adaptability, and exhibits excellent performance and efficiency.
Section 5.1 describes how we organized the experiments, including the settings, datasets, and baselines. Section 5.2 and Section 5.3 present HOOK’s results for sparse and dense tasks on five datasets, with quantitative comparisons in terms of accuracy and efficiency with the previous visual tokenizer and pretrained Patch Embed. The visualization results for the homogeneity of the visual tokens are shown in Section 5.4.
Additionally, we conducted a more in-depth analysis of the phenomena observed in the experiments. Section 5.5 experimentally verifies the simple intuition that under a stronger visual tokenizer, deeper backbone networks are redundant. Section 5.6 discusses whether deeper visual tokenizers perform better. The ablation results and analysis of the two types of attention mechanisms within the OVM are presented in Section 5.7. In Section 5.8, we discuss several limitations of HOOK to provide a comprehensive view of the HOOK method.
@@ 5.1 Setup
Table 1: Results of HOOK on sparse and dense tasks
Tokenizer Sparse Task (Classification) Dense Task (Segmentation)
Number of
Tokens
Params
MACs
NWPU
RS19
NaSC-TG2
Number of
Tokens
Params
MACs
GID5
DGLCC
Patch Embed 196 85.68 16.86 70.30 78.10 86.56 1024 121.47 206.94 67.79 54.57
Quadtree 100 86.02 8.82 70.99 81.05 86.55 400 123.28 158.77 49.18 39.62
VT 8 96.27 2.59 71.95 80.07 84.24 8 133.12 132.83 58.12 45.61
PnP-DETR 54 111.97 8.94 73.53 81.37 85.91 158 148.97 158.82 62.12 50.04
Conv-VGG19 49 105.51 23.78 74.46 83.66 80.60 256 142.62 247.18 53.73 47.44
Conv-ResNet50 49 110.17 8.46 75.17 81.37 86.47 256 147.28 167.15 68.81 48.72
HOOK-32 6 105.51 5.87 71.73 85.29 84.74 8 141.31 147.66 60.98 48.01
HOOK-16 6 99.22 6.80 74.13 84.97 86.94 8 135.01 152.70 73.53 54.08
HOOK-8 6 97.65 10.66 77.21 86.60 87.26 8 133.44 172.87 75.79 54.72
HOOK-4 6 97.25 26.10 77.38 87.58 88.25 8 133.05 253.52 78.81 54.27
@@
A complete transformer-based visual model should consist of three main components: a visual tokenizer, a transformer backbone, and a task head. The visual tokenizer is responsible for converting raw images into token sequences, with the most popular visual tokenizer currently being Patch Embed. The transformer backbone serves as the core of the model, extracting and understanding the semantic information within the token sequences. The task head aligns the output from the backbone with specific downstream tasks. In this paper, we propose a new visual tokenizer called HOOK, which aims to replace Patch Embed to enhance the model’s ability to understand visual data.
@@
To fairly and effectively test the performance of HOOK, unless specified otherwise, we use the same transformer backbone (12-layer, 768-dimensional transformer encoder) and task heads (linear classifier for classification tasks and SegFormer[58] head for segmentation tasks) in all the experiments. Notably, many improvements to visual tokenizers not only change the structure of the visual tokenizer but also introduce new modules or mechanisms into the backbone or task head[28, 30, 38, 42, 59]. To accurately assess the impact of the visual tokenizer on the model’s performance, our baseline for comparison excludes these types of methods and only includes works that focus solely on improving the visual tokenizer.
@@
For the credibility of our experimental results, our first principle in selecting datasets is that they must be publicly available and widely researched. Additionally, we aim to validate that HOOK is image-agnostic and task-agnostic while testing its performance. Therefore, we test the performance of HOOK on both sparse tasks, represented by classification, and dense tasks, represented by segmentation. Specifically, for the classification tasks, we choose NWPU-RESISC45[18], WHU-RS19[19, 20], and NaSC-TG2[21], with all datasets having a fixed image size of 224$\times$224. For the segmentation tasks, we select GID5[22] and DGLCC[23], which have a fixed image size of 512$\times$512.
We use the top-1 accuracy as the evaluation metric for the classification task and the mean intersection over union (mIoU) as the evaluation metric for the semantic segmentation task. We also quantitatively compare the efficiency between different methods, choosing the number of multiply accumulate operations (MACs) as the metric. In all our experiments, the number of MACs refers to the number of multiply accumulate operations during one forwards process of the entire model.
For conciseness, details on dataset preprocessing, experimental hyperparameters, and other specific information are provided in B. Additionally, all executable code, data, and hyperparameters are available in the GitHub repository.
@@ 5.2 Main results
We test the performance of HOOK on two categories of tasks: sparse tasks and dense tasks. The essential difference between the two types of tasks is that sparse tasks require only sparse discriminative information, such as categories, whereas dense tasks require not only discriminative information but also dense spatial information. Representative sparse tasks include classification, reidentification, and sentiment analysis, whereas representative dense tasks include segmentation, detection, and depth estimation.
In addition to the 4$\times$4 pixel seed size, to balance accuracy and efficiency, we also test different HOOK versions with seed sizes of 8$\times$8, 16$\times$16, and 32$\times$32 pixels. As shown in Table 1, these versions are denoted HOOK-4, HOOK-8, HOOK-16, and HOOK-32, respectively.
To accurately evaluate the effects of different visual tokenizers on the overall performance of the model, the baseline for comparison includes only works that optimize the visual tokenizer while keeping the transformer backbone and task head consistent.
@@
Among our selected baseline methods, Quadtree[37] and PnP-DETR[35] are patch-based methods. The former assumes that all patches in the image are unreasonable and uses GradCAM to identify significant regions, implementing the generation of patches with different granularities via quadtrees. The latter reveals that the number of patches is redundant, and a scoring mechanism is used to identify background and foreground regions and merge features of background areas. VT[44] is an object-oriented method that uses a simple convolution operation to assign each pixel in the image to one of several semantic groups, which are eventually mapped to a visual token. Additionally, several studies[27, 35] have suggested that the combination of convolution and a transformer leads to better performance for transformer-based models, so we also tested the performance of two classic convolutional networks (VGG[60] and ResNet[61]) as visual tokenizers.
Table 2: Comparison results of pretrained Patch Embed and HOOK with different ViT architectures
Visual Tokenizer Pretrained Model Publication Arch. Params (M) MACs (G) NWPU RS19
Patch Embed ViT[17] ICLR2023 ViT-Small 21.61 4.25 76.04 82.68
DINO-MC[62] Arxiv2023 21.39 16.73 73.61 78.76
ViT[17] ICLR2021 ViT-Base 52.68 16.86 77.21 84.64
BEiT[63] ICLR2022 85.68 16.86 73.75 82.35
CLIP[64] PMLR2021 85.68 16.86 74.06 83.33
DeiT[65] ICML2021 85.68 16.86 74.89 83.66
MTP[66] JSTARS2024 85.68 16.78 74.27 83.01
CMID[67] TGRS2023 85.68 16.78 72.40 79.08
CSPT[68] RS2022 85.68 16.78 71.82 79.74
ViT[17] ICLR2021 ViT-Large 303.14 59.68 74.94 84.31
BEiT[63] ICLR2022 303.14 59.68 72.98 84.97
CLIP[64] PMLR2021 302.96 77.82 73.28 84.31
SatMAE[69] NIPS2022 303.14 59.38 74.23 83.66
SatMAE-t[69] NIPS2022 303.14 59.38 72.25 79.74
MTP[66] JSTARS2024 303.14 59.38 73.46 84.97
ScaleMAE[70] ICCV2023 303.14 59.38 73.16 82.68
ViT[17] ICLR2021 ViT-Huge 630.49 161.97 75.86 87.58
CLIP[64] PMLR2021 630.49 161.97 73.65 84.31
SAM[55] ICCV2023 630.72 123.57 73.13 83.66
HOOK - - ViT-Small 29.74 25.69 77.32 87.91
ViT-Base 97.25 26.10 77.38 87.58
ViT-Large 318.32 27.42 77.17 87.91
ViT-Huge 650.55 29.41 76.73 86.27
A. Sparse Tasks